OpenStudy (anonymous):

Using the ratio of perfect squares method, what is the square root of 22 rounded to the nearest hundredth?

OpenStudy (mathmate):

The ratio of perfect squares method can tell you that sqrt(22) is between 4 and 5, because 4^2=16<22<25=5^2 Estimating using Newton's method will give you an estimate good to the hundredth and more, as follows. Find the closest integer estimate. 4^2=16=22-6, 5^2=25=22+3, so 5 is closer (but too big). We find the estimate by calculating 5-(5^2-22)/(2*5)=4.70 which is almost good to the hundredth. To get more accuracy, we use 4.7 as the estimate: 4.7-(4.7^2-22)/(2*4.7)=4.6904

OpenStudy (anonymous):

thank you! @mathmate

OpenStudy (mathmate):

You're welcome! :)